OpenStudy (is3535):

A music shop has 5 different brands of guitar strings. The 5 different brands sell for $16, $20, $15, $11, and $13. What is the mean absolute deviation of the prices? Round to the nearest cent, if necessary.

OpenStudy (welshfella):

First work out the mean selling price

OpenStudy (welshfella):

then find the difference between the individual prices and the mean and add them up divide this total by 5

OpenStudy (is3535):

75 divide 5=15

OpenStudy (welshfella):

right now work out 16 - 15 20-15 15-15 15-11 and 15-13

OpenStudy (welshfella):

then add these up and divide by 5 to get the MAD

OpenStudy (is3535):

- or add

OpenStudy (welshfella):

??

OpenStudy (is3535):

15-3 = 12

OpenStudy (welshfella):

why 15 - 3???

OpenStudy (is3535):

now work out 16 - 15 20-15 15-15 15-11 and 15-13

OpenStudy (welshfella):

16 - 15 = 1 20 - 15 = 5 etc

OpenStudy (is3535):

now what

OpenStudy (welshfella):

what do you get after adding the those 5 results?

OpenStudy (is3535):

11

OpenStudy (is3535):

To get mean you need to add up all the numbers and divided it by how many numbers there are so....\[\frac{ 16+20+15+11+13 }{ 5 }\]

OpenStudy (welshfella):

add up the 5 results of 16-15 20 - 15 15-15 15-12 and 15 - 13

OpenStudy (welshfella):

yes - you got the mean already = 15 now we need to find the MAD

OpenStudy (is3535):

15

OpenStudy (is3535):

i think it 9

OpenStudy (welshfella):

yea 15 is the mean then use the mean to work out the MAD by doing those 5 subtractions , adding up the results then divide the total by 5

OpenStudy (welshfella):

gotta go now study all my posts and you'll get it

OpenStudy (is3535):

can you just give me the answer

OpenStudy (is3535):

@Vocaloid

Vocaloid (vocaloid):

well, ok, you just found the mean, which is 15, right?

OpenStudy (is3535):

i got 2.5

Vocaloid (vocaloid):

hm, not quite... check your math again

OpenStudy (is3535):

ok

OpenStudy (is3535):

2.4

Vocaloid (vocaloid):

yup! 2.4 is the right answer

Vocaloid (vocaloid):

@is3535: you can use this website to calculate the mean absolute deviation, so you can check your answers in the future! http://www.alcula.com/calculators/statistics/mean-absolute-deviation/

Vocaloid (vocaloid):

I mean, you still have to calculate it by hand, but you can use the website to see if you did your math right!

OpenStudy (anonymous):

did you learn$\equiv$ and $\pmod{}$yet?